OpenStudy (zuha):

What percent of 94 is 23?  n/94=23/100 23/94=n/100 94/n=23/100 23/94=100/n

OpenStudy (zzr0ck3r):

x94=23, solve for x then make it a percentage

OpenStudy (whpalmer4):

\[23 = x\% \text{ of } 94\]\[x\% = \frac{x}{100}\]\[23 = \frac{x}{100} * 94\] You can rearrange that to be in the same form as one of your answers.

OpenStudy (dwelsher):

simple arithmetic always trumps- 23 is what part of 94? first a fraction and then a division problem

OpenStudy (whpalmer4):

many people asking these questions really have no idea of what a percentage is, and so are likely to construct the wrong arithmetic problem. my hope is always to get the OP to understand the underlying concept(s)...

OpenStudy (zuha):

So what's the answer?

OpenStudy (whpalmer4):

\[23=\frac{n}{100}*94\]What might you do if you were trying to solve that for \(n\)?

OpenStudy (dwelsher):

The first thing to do is to move the 94 to the other side of the equation in the denominator and then cross multipy. another good way to express a fraction(23/94)=(x/100)

OpenStudy (dwelsher):

you have to think about things a little longer to understand this but sooner or later it will sink in

OpenStudy (dwelsher):

what percentage of 8 is 2?, What percentage of 2 is 1? Of course you end up with a decimal. so you have to multiply your answer by 100 to get the correct answer. But as you do math and solve general problems you have to think is this a realistic answer or what am I doing wrong.It better for you to solve it. 23 parts of 94.

OpenStudy (whpalmer4):

and that first step gets you to one of the answers shown above. I completely agree with the opinion about the importance of having a good sense for whether an answer is reasonable! Unfortunately, this problem doesn't give as much scope for exercising it, as it is aimed at testing the student's understanding of the manipulation. But remembering that a percentage is usefully thought of as 'parts of 100', and noticing that 94 is very close to 100, saying 23 is about 23/100 or 23% of 94 gets you a very close approximation! I make my 10 year old son do estimates like this as we walk through stores — "that TV is $499, the Blu-Ray player is $199, estimate the sales tax on the total" etc.